OpenStudy (anonymous):

log (2328) = 2x? Can someone show me how to work this problem

OpenStudy (y2o2):

\[x = {\log(2328) \over 2} = {1\over2 }\log(2328) = \log(\sqrt{2328}) ≈1.6834914879889...\]